[figure: 9.7 cm]

Magnetic properties of NdMn_{2−x}Fe_{x}Ge_{2} (x=0.3, 0.4 and 1.0) compounds.

Ł. Gondek^{a}, M. Duraj^{b}, B. Penc^{a}, N. Stüsser^{c}, Z. Tomkowicz^{a}, A. Szytuła^{a}

aInstutute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków bInstitute of Physics, Technical University of Cracow, Podchora̧żych 1, 30-084 Kraków cBerlin Neutron Scattering Center, Hahn-Meitner Institute, D-14109 Berlin-Wannsee,

Germany

Magnetic properties of the NdMn_{2−x}Fe_{x}Ge_{2} (x=0.3, 0.4 and 1.0) are investigated by AC magnetic measurements and neutron diffraction. In ambient pressure in compound with x=0.3 below 200 K the noncollinear magnetic structure is observed. Above this temperature the magnetic moments transform to the collinear one. For x=0.4 at T=1.5 K the Nd magnetic moments form collinear ferromagnetic structure, while Mn moments noncollinear. With increase of the temperature at T=50 K Nd moment disappears while Mn moments form four different magnetic structures antiferro- (AF) and ferromagnetic (F). The change between AF and F phases is connected with the change of the lattice constants. For x=1.0 the Mn moments form collinear antiferromagnetic structure stable in the temperature range between 1.5 K and T_{N}=203 K. For x=0.3 the external pressure up to 1.5 GPa causes the change of the critical temperature only, while for x=0.4 induces the change of the magnetic properties from ferro- to antiferromagnetic. For x=1.0 the Néel temperature increase with increasing external pressure. The above results indicate that in the NdMn_{2−x}Fe_{x}Ge_{2} system, the increase in the iron content is followed by a change in chemical bond that influences the magnetic properties.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : oral

Corresponding author : A. Szytuła

Address for correspondence : Instutute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków

Email address : szytula@if.uj.edu.pl